If f is integrable on [a, b], the following equation is correct. lim X f (ar) dr f (ari) Aar. where A i 1 Use the given form of the definition to evaluate the integral. and ar a iAr.
Answer #1

\int (2-x^2)dx = 2x-x^3/3 = 2(1-0)-1/3(1^3-0) = 2-1/3 = 6-1/3 = 5/3